OpenStudy (anonymous):

Electrical power of 20 kw is transmitted at 200000 v through a transmission cable having a resistance of 0.1 ohms. A. Calculate the current through the cables. B. Calculate the power wasted in the cables.

OpenStudy (anonymous):

current=2 x 10^6 ampere

OpenStudy (anonymous):

um what formula's do you use? i'm guessing p=vi and v=ir can be used here?

OpenStudy (anonymous):

yea i ussed that formula

OpenStudy (anonymous):

but if i use p=vi for the a part, the answer comes as i=20000/200000 and that would be 0.1 A. but if i use v=ir then the answer comes as 2000000

OpenStudy (anonymous):

hey 2 X10^6=2000000 is same

OpenStudy (anonymous):

yeah. but why is there a difference when two different formula's are used?

OpenStudy (anonymous):

no there isnt for a we will use v=ir for power= p-i^2R or VR

OpenStudy (anonymous):

okay. but why can't we use p=vi for a?

OpenStudy (anonymous):

coz u dont have power supplied in cable

OpenStudy (anonymous):

what do you mean. isn't there an electrical power of 20kw in the cable. so we could re arrange the formula as i=p/v and use it right?

OpenStudy (anonymous):

solve this equation 2x 10^4 (2X10^8-1) so u ll get power

OpenStudy (anonymous):

no u cant a cable alo have it own capability to hold power isntd like in our home 110 volt is upplied but different appliance take diff volt which i used in our home in a regular use it dosent mean whole each appliance will take 110 volt

OpenStudy (anonymous):

yes i know that. but the electrical power supplied through the cable before it reaches our homes is 20000 watts, right? and if the voltage supplied through the same cable is 200000 volts that means we can find the current flowing through that wire or cable, right?

OpenStudy (anonymous):

yea right we can

OpenStudy (anonymous):

but i dunno when i do such qn i alway do such method wat i told u as i m nt dat much good so i can explain but in this wat we have to do is to take out poower and den subtract it from power supplied to get wastage ..

OpenStudy (anonymous):

okay got it. thanks for the reply :)

OpenStudy (anonymous):

yw :)